OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 9x2 − 6x + 1 square feet, what is the length of one side of the room?

OpenStudy (anonymous):

9x^2-6x+1=x^2

OpenStudy (anonymous):

so (9x^2-6x+1)^1/2= length of one side

OpenStudy (anonymous):

thats not an answer choice though

OpenStudy (anonymous):

u gotta factor then

OpenStudy (anonymous):

is 3x-1 an answer?